### v3

# Quantum Aspects of Natural Top Quark Condensation

Abstract
In top quark condensation the Brout-Englert-Higgs (BEH) boson is a $\overline{t}t$ bound state. With a UV completion of a single coloron exchange interaction a recent semiclassical treatment, main ; main2 , gave a novel theory of the BEH boson as an extended object with composite scale $M_{0}\sim 6$ TeV. Presently we obtain the semiclassical theory as an effective action using the source/Legendre transformation techniques of Jackiw, et al. ,\!\!, Jackiw ; CJT , and fermion loop effects in the large-$N_{c}$ limit by deploying an auxiliary field to implement the sum of leading fermion loop diagrams. The theory remains natural at the loop level with fine tuning at the level of a few %, and the effective coupling of the 4-fermion interaction, $\overline{g}_{0}^{2}$, is significantly enhanced by quantum loops over the fundamental coloron coupling, $g_{0}^{2}$. Hence a relatively weaker “topcolor” theory can produce critical coupling in the effective BEH bound state theory.
preprint: Delafield-0820-2025
August 21, 2025
I Introduction
Recently we revisited the idea of “top quark condensation,” i.e., that the Brout-Englert-Higgs (BEH) boson is composed of top + anti-top quarks main ; main2 . Our new approach describes a bound state, consisting of a pair of relativisitic chiral fermions, arising from a single particle exchange of a perturbative massive gluon-like object with mass $M_{0}$ and coupling $g_{0}$ (dubbed a “coloron”). A semiclassical analysis, involving “bilocal fields,” leads to a low-mass composite BEH boson with an extended internal wave-function, $\phi(r)$ main ; main2 . $\phi(r)$ satisfies a Schrödinger-Klein-Gordon (SKG) equation in a Yukawa potential with coupling $g_{0}^{2}$. Close to its critical coupling, $g_{0}^{2}\sim g_{c}^{2}\approx 8\pi^{2}/N_{color}$ the internal wave-function spreads, thus diluting $\phi(0)$. The BEH boson is then an extended object with a small eigenvalue, $\mu^{2}$. We identify $\mu^{2}=-(88)^{2}$ GeV2, the BEH mass in the symmetric phase of the standard model (SM). The composite scale, the mass of the $SU(3)$ octet colorons, is then found to be $M_{0}\sim 6$ TeV. Near critical coupling the low energy theory is approximately scale invariant, in analogy to a second order phase transition in condensed matter physics. This effective scale invariance leads to the spreading of the wave-function $\phi(r)$, and is a universal phenomenon in any bound state system near its critical coupling, where the eigenvalue of the Hamiltonian is small compared to the internal parameters of the theory. Hence scale symmetry is realized dynamically as the custodial symmetry of the small mass BEH bound state.
Remarkably, the static observables, such as the induced top–quark BEH–Yukawa coupling, $g_{Y}$, and quartic coupling, $\lambda$, depend upon $\phi(0)\sim\sqrt{|\mu|/M_{0}}$ and become concordant with experiment main . Moreover, due to a subtlety of the dilution effect, the theory requires fine–tuning at the level of $\sim|\phi(0)|^{2}\sim|\mu|/M_{0}$, rather than the usual $|\mu|^{2}/M_{0}^{2}$. Hence, identifying $\mu^{2}\sim-(88)^{2}$ GeV2 we require fine tuning only at the level of $\sim$ few $\%$. The colorons may be accessible to the LHC, where they would be pair produced and decay to $\overline{t}t$ pairs (also $\overline{b}b$ in extended topcolor models Topcolor ; Topcolor2 ) in boosted jets. This is significantly different than old results obtained in the 1990’s Nambu1 -NSD which were based upon the renormalization group (RG) improved Nambu–Jona-Lasinio model (NJL, NJL ), but did not contain an internal wave-function $\phi(r)$.
Presently we show how the semiclassical theory emerges from the underlying quantum field theory of the third generation quarks with the single coloron exchange interaction. We utilize the formalism developed in the early papers, mainly of Jackiw, et al. ,\!\!, Jackiw CJT , which yield the effective semiclassical action. We also introduce auxiliary fields that facilitate the sum of the leading fermion loop diagrams of order $(\hbar g_{0}^{2}N_{c})^{n}$. The approach can be considered a decomposition of a top quark field into semiclassical + quantum degrees of freedom, $\sim\psi+\sqrt{\hbar}\omega$. Integrating out the quantum fermionic fluctuations, $\omega$, in the presence of fermionic source fields, $\xi_{L,R}$, yields a semiclassical effective “source” action. The action for the semiclassical fields, $\psi_{L,R}(x)$, and a composite bilocal Brout-Englert-Higgs field, $H_{i}(x,y)$, is then generated upon performing a Legendre transformation to remove the sources. The resulting effective action then allows treatment of all semiclassical phenomena in terms of the c-number, chiral spinor fields, $\psi_{L,R}(x)$, and formation of the bound state $H_{i}(x,y)$, reproducing our earlier work (reference main should be considered a companion to the present paper).
Mainly, we find that the sum of $\omega$-loops with coloron coupling, $g_{0}$, generates a renormalized coupling for the 4-fermion interaction, $\overline{g}^{2}_{0}$, where:
$\overline{g}_{0}^{2}=g_{0}^{2}\left(1-\frac{g_{0}^{2}N_{c}}{8\pi^{2}}\right)^{-1}$ (1)
Essentially, the fermion loops act as in the NJL model, which enhances the binding of the $0^{+}$ channel ground state, leading to a form of “critical amplification.” The underlying coloron production, scattering and decay amplitudes are therefore governed by $g_{0}^{2}$, while the binding, which forms the BEH boson, is governed by the enhanced $\overline{g}_{0}^{2}>g_{0}^{2}$. This implies that the $\overline{g}_{0}^{2}$ can be supercritical, while the underlying coloron coupling $g_{0}^{2}$ remains smaller and subcritical. Therefore the effective potential coupling in the $0^{+}$ channel, $\overline{g}_{0}^{2}$, can be super-critical, to produce the low mass BEH bound state and the spontaneous symmetry breaking, in a weaker underlying topcolor model.
The induced top-quark BEH-Yukawa coupling $\overline{g}_{Y}\propto\overline{g}^{2}_{0}$ is also renormalized. Using the experimental value as an input, $\overline{g}_{Y}\approx 1$, we still obtain the composite scale (the coloron mass) $M_{0}\approx 6$ TeV. We also observe how the BEH quartic coupling, ${\lambda}\propto\overline{g}_{0}^{8}\propto\overline{g}_{Y}^{4}$, is generated by the underlying $\omega$-loops. In our present scheme we find numerically that this is remarkably close to the SM result $\lambda\sim 0.25$, which represents a significant improvement over the old NJL model based top condensation that yielded $\lambda\sim 1$.
This also resolves putative issues that arise in the treatment of ref.main . One might be tempted to, e.g., loop the Yukawa interaction to argue that there is a problematic large correction to the BEH mass $\propto-\overline{g}_{Y}^{2}N_{c}M_{0}^{2}/8\pi^{2}$. This would be an additive correction to $\mu^{2}$ and one might conclude that the effective theory is therefore “unnatural.” This would, however, be incorrect. This particular contribution is now already generated by integrating out the underlying $\omega$-fields and it would be double counting to do a fermion loop calculation in the effective semiclassical theory. Moreover, in the underlying theory this appears, not as a mass term, but as a multiplicative and enhancing correction to the effective single coloron exchange interaction and the internal wave-function back-reacts to it. It is, in fact, the source of the critical amplification of $\overline{g}_{0}^{2}$.
We find the Jackiw, et al. ,\!\!, techniques to be powerful, and work particularly well with the auxiliary field fermion loop sum technique introduced here.
II Effective Semiclassical Action
II.1 Sources and Legendre Transformation
We presently derive the effective top quark condensation action for a semiclassical chiral fermion fields $\psi_{iL}(x)$ and $\psi_{R}(x)$, from the underlying quantum field theory, by integrating out third generation, color triplet, quark fields, $(\omega_{iL}(x),\omega_{R}(y))$, where: [^1] Note that the $b_{R}$ quark must also participate in topcolor in an anomaly-free scheme, which can be implemented as in e.g., Topcolor .
$\omega_{iL}=\frac{(1-\gamma^{5})}{2}\biggl{(}\begin{array}[]{c}t\\ b\end{array}\biggr{)},\qquad\omega_{R}=\frac{(1-\gamma^{5})}{2}t,$ (4)
where $i$ is an electroweak $SU(2)$ index. We will use the methods of Jackiw, Jackiw introducing chiral source fields, $\Xi^{i}_{R}$ and $\Xi_{L}$, then performing a Legendre transformation to obtain the effective action for the semiclassical fields $\psi_{L,R}(x)$. We then follow with the method of Cornwall, Jackiw and Tomboulis, (CJT CJT ), in the semiclassical theory and introduce the two-body bound state wave-function source $Q^{i}(x,y)$, together with semiclassical fields. A Legendre transformation then yields the full theory with effective fields $\psi_{L,R}$ and bilocal BEH boson $H_{i}(x,y)$. We will deploy a bilocal auxiliary field, $B(x,y)$, to expedite the sum of fermion loops in a large $N_{c}$ approximation.
Our starting point is the fundamental action at the quantum field theory level of the topcolor model with the single coloron exchange interaction, suitably Fierz rearranged (as given in eq.(28) of ref.main ):
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S(\omega)=\!\int\!d^{4}x\biggl{(}[\overline{\omega}^{i}_{L}(x)\;iD\!\!\!/\!\,\,\omega_{iL}(x)]+[\overline{\omega}_{R}(x)\;iD\!\!\!/\!\,\,\omega_{R}(x)]+g_{0}^{2}\!\int\!d^{4}x\;d^{4}y\;[\overline{\omega}^{i}_{L}(x)\omega_{R}(y)]D_{F}(x-y)[\overline{\omega}_{R}(y)\omega_{iL}(x)],$ (5)
where [][...] denotes color singlet contraction of indices. The coloron propagator in Feynman gauge is:
$iD_{\mu\nu}(x-y)=g_{\mu\nu}D_{F}(x-y);\qquad D_{F}(x-y)=-\int\frac{1}{q^{2}-M_{0}^{2}}e^{iq(x-y)}\frac{d^{4}q}{(2\pi)^{4}}.$ (6)
The interaction can be factorized by introducing a bilocal, isodoublet, auxiliary field $B^{i}(x,y)$ of mass dimension $3$:
$\!\!\!\!\!\!\!\!\!\!S(\omega,B)=\!\int\!d^{4}x\biggl{(}[\overline{\omega}^{i}_{L}(x)\;iD\!\!\!/\!\,\,\omega_{iL}(x)]+[\overline{\omega}_{R}(x)\;iD\!\!\!/\!\,\,\omega_{R}(x)]\biggr{)}$
$\!\!\!\!\!\!\!\!\!\!+\!\int\!d^{4}x\;d^{4}y\;\biggl{(}g_{0}D_{F}(x-y)\left(B^{i\dagger}(x,y)[\overline{\omega}_{R}(y)\omega_{iL}(x)]+[\overline{\omega}^{i}_{L}(x)\omega_{R}(y)]B_{i}(x,y)\right)-[B^{i}{}^{\dagger}(x,y)D_{F}(x-y)B_{i}(x,y)]\biggr{)}.$ (7)
The “equation of motion” of $B_{i}$ is then:
$B_{i}(x,y)=g_{0}[\overline{\omega}_{R}(y)\omega_{iL}(x)];\;\;\qquad B^{i\dagger}(x,y)=g_{0}[\overline{\omega}^{i}_{L}(x)\omega_{R}(y)].$ (8)
Substituting the solution of eq.(8) back into eq.(II.1) yields eq.(5). This is analogous to the treatment of the local NJL model (e.g., as in eqs.(4) and (6) in ref.main ), but differs presently by the bilocal factor, arising from a bilocal coloron exchange amplitude, of $D_{F}(x-y)$. Unlike the NJL model, we do not treat $B_{i}$ as a composite field and we will integrate it out to sum fermion loops.
We can define the fermion kinetic terms and $B_{i}$ interaction terms of eq.(II.1) as a matrix $K(x,y)$:
$S(\Omega,B)=\!\int\!d^{4}x\;d^{4}y\;\biggl{(}[\overline{\Omega}(x)K(x,y)\Omega(y)]-[B^{i}{}^{\dagger}(x,y)D_{F}(x-y)B_{i}(x,y)]\biggr{)},$ (9)
where,
${\Omega}\equiv\begin{pmatrix}\omega_{iL}\\ \omega_{R}\end{pmatrix},\qquad\overline{\Omega}\equiv\begin{pmatrix}\overline{\omega}^{i}_{L}&\overline{\omega}_{R}\end{pmatrix},\qquad K(x,y)=\begin{pmatrix}S_{F}^{-1}(x,y)\;\;&g_{0}D_{F}(x-y)B^{\dagger}(x,y)\\ g_{0}D_{F}(x-y)B(x,y)&S_{F}^{-1}(x,y)\end{pmatrix}$. (10)
Here $S_{F}^{-1}(x,y)$ is the formal inverse of the Feynman propagator, $S_{F}(x,y)=(iD\!\!\!/\!\,\,)^{-1}\delta^{4}(x-y)$. Note that $K$ is Dirac Hermitian, $\overline{K}=\gamma_{0}K^{\dagger}\gamma_{0}=K$.
We follow Jackiw Jackiw and add chiral source terms, $\Xi(x)$, for the fermion fields,
$\!\!\!\!\!\!\!\!\!\!S(\Xi,\Omega,B)=\!\int\!d^{4}x\;d^{4}y\;\biggl{(}[\overline{\Omega}(x)K(x,y)\Omega(y)]-[B^{i}{}^{\dagger}(x,y)D_{F}(x-y)B_{i}(x,y)]\biggr{)}+\!\int\!d^{4}x\biggl{(}\overline{\Xi}(x)\Omega(x)+\overline{\Omega}(x)\Xi(x)\biggr{)},$
$\qquad\qquad\qquad\qquad\qquad\qquad\makebox{where:}\qquad\qquad{\Xi}\equiv\begin{pmatrix}\xi_{iR}\\ \xi_{L}\end{pmatrix},\qquad\overline{\Xi}\equiv\begin{pmatrix}\overline{\xi}^{i}_{R}&\overline{\xi}_{L}\end{pmatrix}$, (11)
(note that the sources have opposite chirality to their corresponding partners, i.e., $\omega_{iL}\leftrightarrow\xi_{iR}$ and $\omega_{R}\leftrightarrow\xi_{L}$). The effective action, $W$, is defined by the path integral:
$W=-i\hbar\ln\!\!\int\!\!D\Omega D\overline{\Omega}\;\exp\biggl{(}\frac{i}{\hbar}S(\Xi,\Omega,B)\biggr{)}.$ (12)
To evaluate $W$ we begin by shifting the integration variables to cancel cross-terms,
$\Omega(x)\rightarrow\Omega(x)-\int\!\!d^{4}y\;K^{-1}(x,y)\Xi(y);\qquad\overline{\Omega}(x)\rightarrow\overline{\Omega}(x)-\int\!\!d^{4}y\;\overline{\Xi}(y)K^{-1}(y,x),$ (13)
and the shifted action becomes:
$S(\Xi,\Omega,B)=\!\int\!d^{4}x\;d^{4}y\;\biggl{(}[\overline{\Omega}(x)K(x,y)\Omega(y)]-\overline{\Xi}(x)K^{-1}(x,y)\Xi(y)-[B^{i}{}^{\dagger}(x,y)D_{F}(x-y)B_{i}(x,y)]\biggr{)}.$ (14)
The path integral is now $W=W(\Omega)+W(\Xi)+W(B)$ where,
$W(\Omega)=-i\hbar\ln\!\!\int\!D\Omega D\overline{\Omega}\;\exp\biggl{(}\frac{i}{\hbar}\!\int\!d^{4}x\;d^{4}y\;\biggl{(}[\overline{\Omega}(x)K(x,y)\Omega(y)]\biggr{)},$
$W(\Xi)=-\!\int\!d^{4}x\;d^{4}y\;\overline{\Xi}(x)K^{-1}(x,y)\Xi(y),\qquad W(B)=-\!\int\!d^{4}x\;d^{4}y\;\biggl{(}[B^{i}{}^{\dagger}(x,y)D_{F}(x-y)B_{i}(x,y)]\biggr{)}.$ (15)
where the quantum loops are contained in $W(\Omega)$. We then define the effective semiclassical field, $\Psi$:
${\Psi}\equiv\begin{pmatrix}\psi_{iL}\\ \psi_{R}\end{pmatrix},\qquad\overline{\Psi}\equiv\begin{pmatrix}\overline{\psi}^{i}_{L}&\overline{\psi}_{R}\end{pmatrix}$ (16)
where,
$\overline{\Psi}(x)=\frac{\delta W(\Xi)}{\delta\Xi(x)}=-\!\int\!\!d^{4}y\;\overline{\Xi}(y)K^{-1}(y,x);\qquad\Psi(x)=\frac{\delta W(\Xi)}{\delta\overline{\Xi}(x)}=-\!\int\!\!d^{4}y\;K^{-1}(x,y)\Xi(y).$ (17)
Hence, inverting these relations,
$\Xi(x)=-\!\int\!\!d^{4}y\;K(x,y)\Psi(y);\qquad\overline{\Xi}(x)=-\!\int\!\!d^{4}y\;\overline{\Psi}(y)K(y,x),$ (18)
$\Xi$ should be regarded as a functional of $\Psi$ by eqs.(18) .
We then perform a Legendre transformation to obtain the effective action, $S(\Psi,B)\equiv W$, swapping $\Xi$ for the semiclassical field, $\psi$, retaining the auxiliary field $B$:
$S(\Psi,B)=W(\Omega)+W(B)+W(\Xi)-\int\!d^{4}x\;\left(\frac{\delta W(\Xi)}{\delta\Xi(x)}\Xi(x)+\overline{\Xi}(x)\frac{\delta W(Xi)}{\delta\overline{\Xi}}\right)$
$=W(\Omega)+W(B)-\!\int\!d^{4}x\;d^{4}y\;\overline{\Xi}(x)K^{-1}(x,y)\Xi(y)-\int\!d^{4}x\;(\overline{\Psi}(x)\Xi(x)+\overline{\Xi}(x)\Psi(x))$
$=W(\Omega)+\int\!d^{4}x\;d^{4}y\;\biggl{(}\overline{\Psi}(x)K(x,y)\Psi(y)-[B^{i}{}^{\dagger}(x,y)D_{F}(x-y)B_{i}(x,y)]\biggr{)}.$ (19)
We obtain the normal-sign kinetic terms and chiralities for the semiclassical fields $\psi_{L,R}$. From eq.(10) we have:
$S(\Psi,B)=W(\Omega)+\!\int\!d^{4}x\biggl{(}[\overline{\psi}_{L}(x)iD\!\!\!/\!\,\,\psi_{L}(x)]+[\overline{\psi}_{R}(x)\;iD\!\!\!/\!\,\,\psi_{R}(x)]\biggr{)}-\int\!d^{4}x\;d^{4}y\;\biggl{(}[B^{i}{}^{\dagger}(x,y)D_{F}(x-y)B_{i}(x,y)]\biggr{)}$
$\qquad\qquad+\!\int\!d^{4}x\;d^{4}y\;\left(g_{0}D_{F}(x-y)(B^{i\dagger}(x,y)[\overline{\psi}_{R}(y)\psi_{iL}(x)]+[\overline{\psi}^{i}_{L}(x)\psi_{R}(y)]B_{i}(x,y)\right).$ (20)
The “equation of motion” of $B_{i}$ is now:
$B_{i}(x,y)=g_{0}[\overline{\Psi}_{R}(y)\Psi_{iL}(x)]+\frac{\delta W(\Omega)}{\delta B^{\dagger i}(x,y)},\qquad B^{i\dagger}(x,y)=g_{0}[\overline{\Psi}^{i}_{L}(x)\Psi_{R}(y)]+\frac{\delta W(\Omega)}{\delta B_{i}(x,y)}.$ (21)
If we ignore the quantum $W(\Omega)$ terms, and insert $B_{i}(x,y)$ back into the new action, we recover the original classical action of eq.(5), but now given in terms of the semiclassical fields $\Psi$:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S(\Psi)=\!\int\!d^{4}x\biggl{(}[\overline{\psi}_{L}(x)iD\!\!\!/\!\,\,\psi_{L}(x)]+[\overline{\psi}_{R}(x)\;iD\!\!\!/\!\,\,\psi_{R}(x)]+g_{0}^{2}\!\int\!d^{4}x\;d^{4}y\;[\overline{\psi}^{i}_{L}(x)\psi_{R}(y)]D_{F}(x-y)[\overline{\psi}_{R}(y)\psi_{iL}(x)].$ (22)
The quantum loop effects are contained in $W(\Omega)$, which we calculate in Section III . We have obtained the semiclassical action for the c-number fields $(\psi_{iL}(x),{\psi}_{R}(y))$.
II.2 Full Effective Action for Bilocal BEH boson via Cornwall-Jackiw-Tomboulis CJT
With the semiclassical theory we now include a source for a bilocal field which will lead to the semiclassical BEH boson field, $H_{i}(x,y)$. Note that, since we have already integrated out the underlying quantum fields, $\omega$, the bound state $H_{i}(x,y)$ that we obtain presently is composed only of the semiclassical modes $\psi_{L,R}$. We can operate at the semiclassical level, following the formalism of Cornwall, Jackiw, and Tomboulis, (CJT) CJT . This is distinct from the NJL model, or the usual applications of CJT, where the bound state is composed in loops of a sum over all of the $\omega$ quantum fields.
We add an additional source term for a complex bilocal field, $Q_{i}(x,y)$, that couples only to a subset of the semiclassical (to-be-bound) modes, $\overline{\psi}_{R}(x)\psi_{iL}(y)]_{b}$, in eqs.(II.1), while the factorization field, $B_{i}$, couples to all fermion pairs, including to-be-free fermion pairs, $\overline{\psi}_{R}(x)\psi_{iL}(y)]_{f}$:
$S(\Psi,B,Q)=W(\Omega)+\!\int_{x}\biggl{(}[\overline{\psi}_{L}iD\!\!\!/\!\,\,\psi_{L}]+[\overline{\psi}_{R}\;iD\!\!\!/\!\,\,\psi_{R}]\biggr{)}$
$+\int_{xy}\;\biggl{(}-[B^{i}{}^{\dagger}D_{F}B_{i}]+\left(g_{0}(B^{i\dagger}+Q^{i\dagger})D_{F}[\overline{\psi}_{R}\psi_{iL}]_{b}+g_{0}B^{i\dagger}D_{F}[\overline{\psi}_{R}\psi_{iL}]_{f}+h.c.\right)\biggr{)},$ (23)
where,
$\int_{x}\biggl{(}[\overline{\psi}_{L}iD\!\!\!/\!\,\,\psi_{L}]+[\overline{\psi}_{R}\;iD\!\!\!/\!\,\,\psi_{R}]\biggr{)}=\int_{x}\biggl{(}[\overline{\psi}_{L}iD\!\!\!/\!\,\,\psi_{L}]_{b}+[\overline{\psi}_{R}\;iD\!\!\!/\!\,\,\psi_{R}]_{b}+[\overline{\psi}_{L}iD\!\!\!/\!\,\,\psi_{L}]_{f}+[\overline{\psi}_{R}\;iD\!\!\!/\!\,\,\psi_{R}]_{f}\biggr{)}.$ (24)
(Here we abbreviate integrals as:
$\int_{u...v}\!\!\!\!=\int\!\!d^{4}u..d^{4}v\;,$ (25)
and we will suppress writing field arguments where obvious, e.g., $B^{i}{}^{\dagger}(x,y)D_{F}(x,y)B_{i}(x,y)\rightarrow B^{i}{}^{\dagger}D_{F}B_{i}$, etc.).
We then have:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{\delta S(\Psi,B,Q)}{\delta Q^{i\dagger}}=g_{0}D_{F}[\overline{\psi}_{R}\psi_{iL}]_{b}\equiv g_{0}\sqrt{N_{c}}M_{0}^{2}D_{F}H_{i};\qquad\frac{\delta S(\Psi,B,Q)}{\delta Q_{i}}=g_{0}D_{F}[\overline{\psi}^{i}_{L}\psi_{R}]_{b}\equiv g_{0}\sqrt{N_{c}}M_{0}^{2}D_{F}H^{i\dagger};$ (26)
where we anticipate the normalization of $H_{i}$, $\overline{\psi}_{R}(x)\psi_{iL}(y)]\rightarrow[\overline{\psi}_{R}(x)\psi_{iL}(y)]_{f}+M_{0}^{2}{\sqrt{N_{c}}}H_{i}(x,y)$, from ref. main . [^2] Only the color singlets bind, hence with color indices $(a,b)$, $\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\overline{\psi}^{a}_{R}(x)\psi_{ibL}(y)\rightarrow\overline{\psi}^{a}_{R}(x)\psi_{ibL}(y)_{f}+M_{0}^{2}\frac{\delta{{}^{a}_{b}}}{\sqrt{N_{c}}}H_{i}(x,y).$ (27) which yields the $\sqrt{N_{c}}$ factors in eq.(26). Note that in main we used $M^{2}=\epsilon M_{0}^{2}$ as an arbitrary normalizing mass scale for $H$, but we then showed that $\epsilon\rightarrow 1$ in the critical coupling limit. We are presently abbreviating the discussion and assume $\epsilon=1$ at the outset.
We perform the Legendre transformation, where $Q$ is now viewed as a functional of $H$ by eq.(26), following CJT :
$S(\Psi,B,H)=S(\Psi,B,Q)-\int_{xy}\biggl{(}Q^{\dagger}\frac{\delta S}{\delta Q^{\dagger}}+Q\frac{\delta S}{\delta Q}\biggr{)},$ (28)
hence,
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S(\psi,B,H)=W(\Omega)+\!\int_{x}\biggl{(}[\overline{\psi}_{L}iD\!\!\!/\!\,\,\psi_{L}]+[\overline{\psi}_{R}\;iD\!\!\!/\!\,\,\psi_{R}]\biggr{)}+\int_{xy}\;\biggl{(}-[B^{i}{}^{\dagger}D_{F}B_{i}]+g_{0}(B^{i\dagger}D_{F}C_{i}+h.c.)\biggr{)},$ (29)
where we define,
$C_{i}(x,y)=\sqrt{N_{c}}M_{0}^{2}H_{i}(x,y)+[\overline{\psi}_{R}(y)\psi_{iL}(x)]_{f}.$ (30)
If we ignore the quantum corrections ($\omega$-loops), the “equation of motion” of $B_{i}$ becomes:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!B_{i}(x,y)=g_{0}C_{i}(x,y),\qquad B^{i\dagger}(x,y)=g_{0}C^{i\dagger}(x,y).$ (31)
Substituting $B_{i}$ into eq.(29) we thus obtain the semiclassical result,
$S(\Psi,H)=\!\int_{x}\!\biggl{(}[\overline{\psi}_{L}(x)iD\!\!\!/\!\,\,\psi_{L}(x)]_{b}+[\overline{\psi}_{R}(x)\;iD\!\!\!/\!\,\,\psi_{R}(x)]_{b}+[\overline{\psi}_{L}(x)iD\!\!\!/\!\,\,\psi_{L}(x)]_{f}+[\overline{\psi}_{R}(x)\;iD\!\!\!/\!\,\,\psi_{R}(x)]_{f}\biggr{)}$
$+g_{0}^{2}\!\int_{xy}\biggl{(}N_{c}M_{0}^{4}H^{\dagger}(x,y)D_{F}(x-y)H(x,y)+\;\sqrt{N_{c}}M_{0}^{2}(H^{i\dagger}(x,y)D_{F}(x-y)[\overline{\psi}_{R}(y)\psi_{iL}(x)]_{f}+h.c.)$
$\qquad\qquad+[\overline{\psi}^{i}_{L}(x)\psi_{R}(y)]_{f}D_{F}(x-y)[\overline{\psi}_{R}(y)\psi_{iL}(x)]_{f}\biggr{)}.$ (32)
We have therefore obtained results identical to the starting point of our previous theory, eq.(54), in ref.main . Note the BCS-like enhancement factor of $N_{c}$ in the $H^{\dagger}D_{F}H$ interaction term, where $N_{c}$ is the analogue of the number of Cooper pairs in a BCS state BCS .
II.3 Kinetic Terms and Factorization Ansatz
The “to-be-bound” fermion kinetic terms remain in the action $S^{\prime}$:
$S_{KTb}=\!\int_{x}\!\biggl{(}[\overline{\psi}^{i}_{L}(x)iD\!\!\!/\!\,\,\psi_{iL}(x)]_{b}+[\overline{\psi}_{R}(x)\;iD\!\!\!/\!\,\,\psi_{R}(x)]_{b}\biggr{)},$ (33)
These have to be replaced by the kinetic term of the bound state field $H_{i}(x,y)$. From eq.(26), in the free field limit (small $g_{0}$), the kinetic terms $S_{KTb}$ imply the equations of motion. Technically with $H\equiv H_{i}\frac{\tau^{i}}{2}$ these are best written as commutators,
$[D_{y},[D_{y},H(x,y)]]=0\qquad[D^{\dagger}_{x},[D^{\dagger}_{x},H(x,y)]]=0.$ (34)
Omitting complications from the gauge fields these follow from the square of the Dirac equation, e.g. $\overline{\psi}_{L}(y)[\partial\!\!\!/\!\,\,,[\partial\!\!\!/\!\,\,,\psi_{iR}(x)]]_{b}=\partial_{x}^{2}[\overline{\psi}_{L}(y)\psi_{iR}(x)]]_{b}=0$, etc. A free field Lorentz invariant action yields the equations of motion by variation:
$\!\!\!\!\!\!\!\!\!\!S_{KTb}\rightarrow S_{KTH}=M_{0}^{4}\int\!d^{4}x\;d^{4}y\;\biggl{(}Z|D_{R}^{\dagger}H(x,y)|^{2}+Z|D_{L}H(x,y)|^{2}\biggr{)},$ (35)
where the covariant derivatives are as defined in the NJL model:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!D_{L\mu}=\frac{\partial}{\partial y_{\mu}}-ig_{2}W_{\mu}^{A}(y)\frac{\tau^{A}}{2}-ig_{1}W_{\mu}^{0}(y)\frac{Y_{L}}{2};\qquad D^{\dagger}_{R\mu}=\frac{\partial}{\partial x_{\mu}}+ig_{1}W_{\mu}^{0}(x)\frac{Y_{R}}{2},$ (36)
where $W^{A}_{\mu}$ ($W^{0}_{\mu}$) denotes the $SU(2)$ $U(1)_{Y}$ gauge fields and the weak hypercharges are $Y_{L}=1/3$, $Y_{tR}=4/3$, (and the presently unused $Y_{bR}=-2/3$), and we will require $Y_{H}=-1$ for the BEH boson of the SM; the electric charges are as usual: $Q=I_{3}+\frac{Y}{2}$. Note that $D_{L}$ ($D^{\dagger}_{R}$) acts at coordinate $y$ ($x$), and $D^{\dagger}_{R}$ acts on $\overline{t}_{R}$, hence the sign flip in the gauge field terms (the derivative $D^{\dagger}$ acts in the forward direction as we have written the kinetic term eq.35).
We now pass to barycentric coordinates,333 Note: the Jacobian was misquoted in the preprint version of main but corrected in the published version.
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!X^{\mu}=\frac{x^{\mu}+y^{\mu}}{2},\;\;\;r^{\mu}=\frac{x^{\mu}-y^{\mu}}{2},\;\;\;\partial_{x}=\frac{1}{2}(\partial_{X}+\partial_{r}),\;\;\;\partial_{y}=\frac{1}{2}(\partial_{X}-\partial_{r})\;\;\;\makebox{Jacobian:}\;\;\;\left|\frac{\partial(X,\rho)}{\partial(x,y)}\right|\equiv J^{-1}=2^{-4}.$ (37)
To work in the rest frame as in main we follow Yukawa, Yukawa , and for the ground state we consider a factorized ansatz for $H_{i}(X,r)$: [^4] Note that we have not specified that $H(x,y)$ or $Q(x,y)$ is factorized in $(x,y)$, as $\sim\overline{\psi}_{R}(x)\psi_{L}(y)$, hence we are permitted to make the ansatz $H_{i}(X)\phi(r)$. An $(X,r)$ factorized form is a mixture of $(x,y)$ factorized basis functions, e.g., $H(X)\phi(r)\sim\sum_{ij}a_{ij}\overline{\psi}_{Ri}(x)\psi_{Lj}(y)$, i.e., the boundstate is an entangled combination of free scattering states.
$H_{i}(x,y)=H_{i}(X+r,X-r)\equiv\hat{H}_{i}(X,r)\qquad\text{hence,}\qquad\sqrt{J/2}\;\hat{H}_{i}(X,r)=H_{i}(X)\phi(r).$ (38)
We can introduce Wilson lines as an approximation to “pull-back” the gauge couplings from $(x,y)$ to the center $X$. This is done by field redefinitions (as discussed in main ):
$H_{i}(x,y)\rightarrow W^{\dagger}_{R}(X,x)W_{L}(X,y)\;\sqrt{2/J}\;H_{i}(X)\phi(r),$ (39)
where $H_{i}(X)$ will have mass-dimension 1, and and $\phi$ will be dimensionless. Note that $H_{i}(X)$, is distinguished from $H_{i}(x,y)$ by having dependence only upon $X$, and will become the effective BEH boson. The Wilson lines are not necessary. In the pure gauge limit these are simply gauge transformations on the $H_{i}$ field. We introduce the Wilson lines as a conceptual tool to consolidate the electroweak charges at $X$, though the gauge charges are effectively consolidated at $X$ in the low energy approximation. Moreover, the Wilson lines are renormalizable and act as sources (charges) only at their end-points. However, they do radiate, acting as currents.
With the pull-back localization of the BEH electroweak charges, the $W$ and $Z$ masses are generated in the usual way under spontaneous symmetry breaking. [^5] We have neglected the gluon Wilson line that extends from $\vec{r}$ to $\vec{r}$, but of course, produces no color charge of the BEH boson. We have the Lorentz invariant kinetic term action:
$S_{KTH}\rightarrow\!\!\int\!\!d^{4}Xd^{4}r\;ZM_{0}^{4}\biggl{(}|D_{H}H(X)|^{2}|\phi(r)|^{2}+|H(X)|^{2}|\partial_{r^{\mu}}\phi(r)|^{2}\biggr{)},$ (40)
where the covariant derivative is:
$D_{H\mu}=\frac{\partial}{\partial X^{\mu}}-ig_{2}W^{A}_{\mu}(X)\frac{\tau^{A}}{2}-ig_{1}W^{0}_{\mu}(X)\frac{Y_{H}}{2}.$ (41)
The internal wave-function $\phi(r)$ is now a complex scalar that carries no gauge charges, and $Y_{H}=-Y_{tR}+Y_{L}=-1$, apropos the BEH field. (see Appendix A of main for the algebra of the Wilson line pullback).
II.4 Removing Relative Time
Consider the kinematics of a pair of massless particles of 4-momenta $p_{1}$ and $p_{2}$. We have $p_{1}^{2}=p_{2}^{2}=0$. and two-body plane waves, $H(x,y)\sim\exp(ip_{1}x+ip_{2}y)$. We pass to the total momentum $P=(p_{1}+p_{2})$ and relative momentum $Q=(p_{1}-p_{2})$, and the plane waves become $\exp(iPX+iQr)$. Note that $P_{\mu}Q^{\mu}=p_{1}^{2}-p_{2}^{2}=0$. This implies in the rest (barycentric) frame, $P=(P_{0},0)$ and $Q=(0,\vec{Q})$, and the free equation of motion, $P_{0}^{2}-\vec{Q}^{\;2}=0$. Hence, in the rest frame the dependence upon $\vec{X}$ (associated with $\vec{P}$) and, in particular the relative time, $r^{0}$ (associated with $Q^{0}$), drops out. If the particles are constituents of a bound state then this is the rest frame of the composite particle. In any other frame there is always a boosted $r^{0}{}^{\prime}$ that is unphysical.
Bilocal fields have spurious dependence upon $r^{0}$ in the rest frame (or its boosted $r^{0}{}^{\prime}$ in any other frame). This is an unphysical and unwanted degree of freedom. It is analogous to the situation in a gauge theory where $\partial_{\mu}A^{\mu}$ is unphysical. [^6] In a gauge theory, such as QED, one may think this is relevant only to define the path integral. However this is an issue in classical electrodynamics, as well, and leads to the “transverse current” in general gauges Jackson To use the bilocal formalism, $r^{0}{}^{\prime}$ must be projected out with Lorentz invariant constraints, analogous to gauge fixing. We have discussed this in greater detail in the present action formalism elsewhere main ; main2 .
In the action of eq.(40) we included a normalization factor $Z$ as in main . This is natural in mapping from the fermionic kinetic terms to the $\hat{H}_{i}$ kinetic terms, but it serves the role of allowing us to remove relative time. In the barycentric (center-of-mass) frame we assume an ansatz for $\phi(r)$ ($H_{i}(X))$ that has no dependence upon the relative time $r^{0}$ ($\vec{X}$). The field $H_{i}(X)$ will have the conventional volume normalization of a mass-dimension 1, complex field. The field $\phi(\vec{r})$ however, will describe the bound state and must be a compact (normalizable) field.
A canonical normalization of the $H_{i}(X)$ kinetic term in any frame (which leads to a normalized Noether current, such as $iH^{\dagger}\overleftrightarrow{\partial}_{\!\!\!X}H$) dictates a Lorentz invariant normalization constraint:
$1=\int d^{4}r\;ZM_{0}^{4}\;|\phi(r)|^{2}.$ (42)
Since the relative time disappears kinematically in the rest frame, then $\phi(r)\rightarrow\phi(\vec{r})$ becomes a static field that has no dependence upon $r^{0}$. We then define the compact normalization for $\phi(\vec{r})$ in the rest frame as:
$1=M_{0}^{3}\int d^{3}r\;|\phi(\vec{r})|^{2},\qquad\text{which implies,}\qquad\int ZM_{0}\;dr^{0}=1,\qquad\text{(see below).}$ (43)
These conditions can be specified in a manifestly covariant form using the unit vector $\omega_{\mu}=P_{\mu}/\sqrt{P^{2}}$, and e.g., $r^{0}=r_{\mu}\omega^{\mu}$, etc., as in Appendix C of main . Note that $\phi(\vec{r})$ is dimensionless with eq.(43).
This can be viewed as a “prenormalization” to define the free bilocal field and it’s Noether charges main2 , before turning on the interaction. The factor $Z$ is not absorbed into the fields after turning on the interaction since we don’t want the coupling constant, $g_{0}^{2}$, to be renormalized at the classical level. Indeed, $Z$ can be viewed as an operator, $Z\sim\delta(M_{0}r^{0})$. We remark that all field theories, even local ones, are implicitly defined with a “canonical” prenormalization that defines the semiclassical action after which we turn on the interaction. Then, at loop level, we are compelled to renormalize the fields to preserve the canonicality which, in turn, renormalizes the coupling constants.
We see that $S_{KTH}$ becomes,
$S_{KTH}\rightarrow\!\!\int\!\!d^{4}X\biggl{(}|D_{X}H(X)|^{2}+|H(X)|^{2}M_{0}^{3}\!\!\int\!\!d^{3}r\;\left(-|\partial_{\vec{r}}\phi(\vec{r})|^{2}\right)\biggr{)}.$ (44)
We then turn on the interaction with coupling constant $g_{0}$ of eq.(II.2), and the semiclassical action becomes,
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S_{0}(H)=\!\!\int\!\!d^{4}X\biggl{(}|D_{X}H(X)|^{2}+|H(X)|^{2}M_{0}^{3}\!\!\int\!\!d^{3}r\;\left(-|\partial_{\vec{r}}\phi(\vec{r})|^{2}\right)+|H(X)|^{2}\!\!\int\!\!dr^{0}d^{3}r\left(M_{0}^{4}\;{2g_{0}^{2}N_{c}}D_{F}(2r^{\mu})|\phi(\vec{r})\;|^{2}\right)\!\biggr{)}.$ (45)
Note that there is no $Z$ factor in the interaction term, and $|\partial_{r^{\mu}}\phi|^{2}=|\partial_{r^{0}}\phi|^{2}-|\partial_{\vec{r}}\phi|^{2}\rightarrow-|\partial_{\vec{r}}\phi|^{2}$. If we were to define $z\equiv\int ZMdr^{0}$ , then we can rescale $\phi\rightarrow\phi/\sqrt{z}$ to have the compact normalization of eq.(43), but then the coupling constant becomes $g_{0}^{2}/z$. However, we want the theory to have at the classical level the same coupling $g_{0}^{2}$ as defined by the coloron exchange, e.g., we want to match the low energy scattering matrix elements that we would obtain in the conventional topcolor field theory to the bilocal theory. Hence $z=1$ is a matching condition of the underlying topcolor interaction to the bilocal theory.
We therefore integrate over $r^{0}$ in the interaction term :
$\!\!\!\!\!\!\!\!\!\!\!\!\!\int\!\!dr^{0}D_{F}(2r)=-\!\!\int\!\!dr^{0}\frac{d^{4}q}{(2\pi)^{4}}\frac{1}{q^{2}-M_{0}^{2}}e^{2iq_{\mu}r^{\mu}}=\frac{1}{2}\int\!\!\frac{d^{3}q}{(2\pi)^{3}}\frac{1}{\vec{q}{\;}^{2}+M_{0}^{2}}e^{2iq_{\mu}r^{\mu}}=-\frac{1}{2}V_{0}(2|\vec{r}|)=\frac{e^{-2M_{0}|\vec{r}|}}{16\pi|\vec{r}|}.$ (46)
The $\vec{q}$ momentum integral yields the familiar Yukawa potential (where $2|\vec{r}|$ is the separation of the particles). The action then becomes,
$\!\!\!\!\!\!\!\!\!\!S=\!\!\int d^{4}X\biggl{(}|D_{X}H(X)|^{2}+|H(X)|^{2}M_{0}^{3}\!\!\int\!\!d^{3}r\left(-|\partial_{\vec{r}}\phi(\vec{r})|^{2}+{g_{0}^{2}N_{c}M_{0}}\frac{e^{-2M_{0}|\vec{r}|}}{8\pi|\vec{r}|}|\phi(\vec{r})|^{2}\right)\biggr{)}.$ (47)
While not “manifestly’ so, this is “implicitly” Lorentz invariant as discussed in Appendix C of main . Note that, in the limit of suppressing the $\vec{q}{\;}^{2}$ in the denominators of the integrands of eq.(46), we obtain the large $M_{0}^{2}$ limit of the potential (using $J=2^{4}$, and $\delta^{3}(\vec{r})=(4\pi r^{2})^{-1}\delta(r)$):
$V_{0}(2r)\rightarrow-\frac{1}{M^{2}_{0}}\delta^{3}(2\vec{r})=-\frac{1}{8M^{2}_{0}}\delta^{3}(\vec{r})\qquad D_{F}(2r)\rightarrow\frac{1}{M^{2}_{0}}\delta^{4}(2r^{\mu})=\frac{1}{JM^{2}_{0}}\delta^{4}(r^{\mu}).$ (48)
This recovers the NJL model potential.
The manipulation leading to eq.(47) may seem to be inconsistent, but it is a short-cut for a more formal limiting procedure. In a manifestly Lorentz invariant form of the internal wave-function action of eq.(45),
$\!\!\int\!\!d^{4}r\;\biggl{(}Z|\partial_{r^{\mu}}\phi(r^{\mu})|^{2}+\;{2g_{0}^{2}N_{c}}D_{F}(2r^{\mu})|\phi(r^{\mu})\;|^{2}\!\biggr{)},$ (49)
we want to take $\phi(r^{\mu})\rightarrow\phi(\vec{r})$, which is independent of $r^{0}$. However this appears to be inconsistent with the interaction term with the factor $D(2r^{\mu})$ which has nontrivial dependence upon $r^{0}$. The problem arises from having integrated out non-static colorons to produce $D_{F}(2r^{\mu})$. However to obtain a Hamiltonian (such as in functional Schrödinger picture) we go to static field configurations. To see this, consider a toy constituent, complex scalar, field theory:
$S=\!\!\int\!\!d^{4}x\biggl{(}(|\partial\varphi_{1}(x)|^{2}+|\partial\varphi_{2}(x)|^{2}-gM_{0}(\varphi_{1}(x)\varphi_{2}(x)A(r)+h.c.)+|\partial A(x)|^{2}-M_{0}^{2}|A(x)|^{.}\biggr{)}.$ (50)
The Hamiltonian for this theory takes the form
$H=\!\!\int\!\!d^{3}x\;\biggl{(}|\pi_{1}(\vec{x})|^{2}+|\pi_{2}(\vec{x})|^{2}+|\nabla\varphi_{1}(\vec{x})|^{2}+|\nabla\varphi_{2}(\vec{x})|^{2}+gM_{0}(\varphi_{1}(\vec{x})\varphi_{2}(\vec{x})A(\vec{x})+h.c.)$
$\qquad\qquad+|\pi_{A}(\vec{x})|^{2}+|\nabla A(\vec{x})|^{2}+M_{0}^{2}|A(\vec{x})|^{2}\biggr{)},$ (51)
where the canonical momenta are $\pi_{i}=\partial_{0}\varphi_{i}$ and $\pi_{A}=\partial_{0}A$. We are interested in the static limit of this theory so we impose the constraints, $\pi_{i}=\pi_{A}=0$, and there is then no Poisson bracket (hence no Dirac bracket). The static equation for $A(\vec{x})$ is, $\nabla^{2}A-M_{0}^{2}A=gM_{0}\varphi_{1}\varphi_{2}$, hence we can integrate out $A$, and write,
$H=\!\!\int\!\!d^{3}xd^{3}y\;\biggl{(}ZM_{0}|\nabla_{x}\varphi_{1}(x)|^{2}|\varphi_{2}(y)|^{2}+ZM_{0}|\nabla_{y}\varphi_{2}(y)|^{2}|\varphi_{1}(x)|^{2}-g^{2}M^{2}_{0}|\varphi_{1}(x)\phi_{2}(y)|^{2}\Delta(x-y)\biggr{)},$ (52)
where $1=ZM_{0}\int d^{3}y|\varphi_{i}(y)|^{2}$, and $\Delta(x-y)=V_{0}(\vec{x}-\vec{y})/2$. Now, replace $\varphi_{1}(x)\varphi_{2}(y)\rightarrow M_{0}\Phi(x,y)$ to obtain,
$H=\!\!\int\!\!M_{0}^{3}d^{3}xd^{3}y\;\biggl{(}Z|\nabla_{x}\Phi(x,y)|^{2}+Z|\nabla_{y}\Phi(x,y)|^{2}-\frac{1}{2}g^{2}M_{0}|\Phi(x,y)|^{2}V_{0}(x-y)\biggr{)}.$ (53)
We go to the barycentric coordinates, and factorize $\Phi(x,y)=\sqrt{2/J}\;\chi(X)\phi(\vec{r})$, [^7] A c-number product $\varphi_{1}(\vec{x})\varphi_{2}(\vec{y})$ cannot generally be rewritten in the factorized form, but quantum fields can lead to states that are so factorized, $\langle 0|\varphi_{1}(\vec{x})\varphi_{2}(\vec{y})|S\rangle\sim\chi(X)\phi(\vec{r})$.
$H=\!\!\int\!\!M_{0}^{3}d^{3}Xd^{3}r\;\biggl{(}Z|\nabla_{X}\chi(\vec{X})|^{2}|\phi(\vec{r})|^{2}+Z|\chi(\vec{X})|^{2}|\nabla_{\vec{r}}\phi(\vec{r})|^{2}-g_{0}^{2}N_{c}M_{0}|\chi(\vec{X})|^{2}|\phi(\vec{r})|^{2}V_{0}(2\vec{r})\biggr{)},$ (54)
with $g^{2}=g_{0}^{2}N_{c}$. We can do the same construction with the full coloron theory. The above example displays the origin of $Z$ and shows that there is no inconsistency in the final result of eq.(47) obtained by the short-cut which yields the single particle exchange interaction as it would be generated in the static frame. This is the correct procedure leading to a Hamiltonian for the functional Schrödinger equation of the field theory with the bound state (see the discussion in CJT of Kuti’s method). To recover the action formalism we would do a Legendre transformation from $H$ back to the Lagrangian $L$, restoring the time derivative on $\chi(X)$ and integrating $L$ over $X^{0}$.
We emphasize that we can often use the NJL limit of the potential, but the bilocal field remains. For example, in the Dirac $\delta$-function potential of $1+1$ quantum mechanics, the wave-function $\phi(r)$ is extended in space even though the potential is pointlike main . In $1+3$ dimensions the Dirac $\delta$-function potential theory is actually ambiguous, and depends upon the limiting procedure used to define it. This is the analogue of what we are facing presently, but the Yukawa potential affords a well defined limiting procedure and unambiguous solutions.
Including the Yukawa interaction and a loop generated quartic term we obtain the effective action for the composite BEH field,
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S=\!\!\int\!\!d^{4}X\;\biggl{(}|D_{H}H(X)|^{2}+|H(X)|^{2}\;M_{0}^{3}\!\!\int\!\!d^{3}r\;\biggl{(}-|\partial_{\vec{r}}\phi(r)|^{2}+{g_{0}^{2}N_{c}M_{0}}\frac{e^{-2M_{0}|\vec{r}|}}{8\pi|\vec{r}|}|\phi(\vec{r})|^{2}\biggr{)}$
$\qquad\qquad-\frac{\lambda}{2}(H^{\dagger}(X)H(X))^{2}-g_{Y}\left([\overline{\psi}^{i}_{L}(X)t_{R}(X)]_{f}H_{i}(X)+h.c.\right)\biggr{)}.$ (55)
The internal field $\phi(r)$ is “nested” within the action for a conventional pointlike BEH boson, $H(X)$. The static $\phi$ field has a Hamiltonian:
${\cal{M}}=M_{0}^{3}\!\!\int\!\!d^{3}r\;\biggl{(}|\partial_{\vec{r}}\phi(r)|^{2}-{g_{0}^{2}N_{c}M_{0}}\frac{e^{-2M_{0}|\vec{r}|}}{8\pi|\vec{r}|}|\phi(\vec{r})|^{2}\biggr{)}.$ (56)
[…]
$-\nabla^{2}\phi-g_{0}^{2}N_{c}M_{0}\frac{e^{-2M_{0}|\vec{r}|}}{8\pi|\vec{r}|}\phi(r)=\mu^{2}\phi.\qquad\text{where,}\qquad\nabla^{2}=\frac{\partial^{2}}{\partial r^{2}}+\frac{2}{r}\frac{\partial}{\partial r}.$ (57)
We find that the SKG equation has a critical coupling, $g^{2}_{c}N_{c}/8\pi^{2}=1.06940$, for which $\mu^{2}=0$. Remarkably, this is very close to the quantum NJL critical coupling $g^{2}_{NJLc}N_{c}/8\pi^{2}=1.00$ (see Section IV. (B) of main ). When $g^{2}_{0}>g^{2}_{c}$ the eigenvalue $\mu^{2}$ becomes negative, $\rightarrow-|\mu|^{2}$. In such a solution the action for $H(X)$ then becomes the familiar,
$\!\!\!\!\!\!\!\!\!\!\!\!\!S=\!\!\int d^{4}X\biggl{(}|D_{H}H(X)|^{2}+|\mu|^{2}|H(X)|^{2}-\frac{\lambda}{2}(H^{\dagger}(X)H(X))-g_{Y}\left([\overline{\psi}_{L}(X)t_{R}(X)]_{f}H(X)+h.c.\right)\biggr{)},$ (58)
with the “sombrero potential”:
$-|\mu|^{2}(H^{\dagger}(X)H(X))+\frac{\lambda}{2}(H^{\dagger}(X)H(X))^{2}.$ (59)
As discussed in detail in main , the solution of the Schrödinger-Klein-Gordon SKG equation for $\phi(r)$ indeed extends to large distances, $\phi(r)\sim e^{-|\mu|r}/r$ where $|\mu|<\!\!M_{0}$ near critical coupling. This dilutes the value of $\phi(0)\sim\sqrt{|\mu|/M_{0}}$. We find that the Yukawa coupling $g_{Y}\propto\phi(0)$ and $\lambda\propto g_{Y}^{4}\propto|\phi(0)|^{4}$. Moreover, inputting the known value of the Lagrangian mass of the BEH boson in the symmetric phase, $|\mu|^{2}=-(88)^{2}$ GeV2, we find the composite scale (coloron mass) $M_{0}\approx 6$ TeV. The quartic coupling, $\lambda$, will be determined at loop level and we find remarkable agreement with the SM as discussed below.
The degree of fine-tuning of the theory is also suppressed by $\phi(0)$ in a subtle way. Rather than the naive result one would expect from the NJL model, $\delta g_{0}^{2}/g_{c}^{2}\sim|\mu|^{2}/M_{0}^{2}\sim 10^{-4}$, we now obtain a linear relation: $\delta g_{0}^{2}/g_{c}^{2}\sim|\mu|/M_{0}\sim 1\%$. To obtain these results we need the precise relationship between $g_{Y}$ and $\phi(0)$ which follows from the induced Yukawa coupling of the bound state to free fermions.
II.5 The Induced Bound State Yukawa Interaction
We see in eq.(II.2) that the Yukawa interaction of the bound state with the free scattering state fermions is now induced from the $\sqrt{N_{c}}$ term:
$\!\!\!\!\!\!\!\!\!\!g^{2}_{0}\sqrt{2N_{c}J}M_{0}^{2}\int\!\!d^{4}Xd^{4}r\;[\overline{\psi}^{i}_{L}(X\!+\!r)\psi_{R}(X\!-\!r)]_{f}D_{F}(2r)\;H_{i}(X)\phi(\vec{r}){+h.c.}.$ (60)
Consider the pointlike limit of the potential of eq.(48), $D_{F}(2r)\rightarrow(JM_{0}^{2})^{-1}\delta^{4}(r)$:
$\rightarrow g^{2}_{0}\sqrt{2N_{c}J}\int\!\!d^{4}X\;[\overline{\psi}^{i}_{L}(X)\psi_{R}(X)]_{f}H_{i}(X)\phi(0){+h.c.}.$ (61)
We therefore see that the induced Yukawa coupling to the field $H(x)$ is:
$g_{Y}=g_{0}^{2}\sqrt{2N_{c}/J}\;\phi(0).$ (62)
Here the behavior of $\phi(0)$ is a suppression of $g_{Y}$ that is a power-law, $\sim\sqrt{|\mu|/M_{0}}$ for small $\mu$, near the critical coupling. This power-law behavior is significantly different than the slow RG (logarithmc) evolution in the old NJL-based top condensation model, which is why we now obtain $M_{0}\sim 6$ TeV and reduced fine-tuning $\sim|\phi(0)|^{2}\sim$ few $\%$ (rather than $M_{0}\sim 10^{15}$ GeV and extreme fine-tuning $\sim 10^{-26}$).
III Quantum Fermion Loop Corrections to the Semiclassical Theory
We recall the results for the fermion loops that arise in the pointlike NJL model with Yukawa interaction:
$\int_{x}\biggl{(}g_{0}[\overline{\psi}^{i}_{L}(x)\psi_{R}(x)]H_{i}(x)+h.c.\biggr{)}.$ (63)
This leads to the loop-induced, ${\cal{O}}(\hbar)$, terms in an effective action (ignoring classical terms):
$\!\!\!\!\!\!\!\!\!\!S_{eff}=\!\int\!d^{4}x\biggl{(}\widetilde{Z}DH^{\dagger}DH+M^{2}H^{\dagger}H-\frac{\lambda}{2}(H^{\dagger}H)^{2}\biggr{)},$ (64)
where,
$\!\!\!\!\!\!\!\!\!\!\widetilde{Z}=\frac{g_{0}^{2}N_{c}}{8\pi^{2}}\ln(M_{0}/m),\qquad M^{2}=\!\frac{g_{0}^{2}N_{c}}{8\pi^{2}}M_{0}^{2},\qquad\lambda=\frac{g_{0}^{4}N_{c}}{4\pi^{2}}\ln(M_{0}/m).$ (65)
The sign of the loop-induced $M^{2}$ term in the action is positive, representing an attractive negative term (tachyonic mass) in the effective potential, while the quartic term is negative in the action, hence, repulsive in the potential. In the NJL model, with classical bare mass $M_{0}^{2}$, the resulting physical mass of the pointlike bound state is $\mu^{2}=\widetilde{Z}^{-1}(M_{0}^{2}-M^{2})$, which defines the NJL critical coupling, $1=g_{0}^{2}N_{c}/8\pi^{2}$. Hence $M^{2}$ is the analogue of “binding energy” associated with the bound state.
This is a hint that the quantum loop effects will enhance the binding of the interaction in the semiclassical bilocal theory. However, it is important to allow the internal wave-function to respond to the attractive $M^{2}$ term, which now becomes part of the potential. Moreover, the NJL limit is a $\delta$-function approximation, and we expect in the bilocal theory a softer contribution to the short distance potential than the mass $M^{2}$. Indeed, in the the loop calculation the term $M_{0}^{2}\delta^{4}(2r)$ is indistinguishable from the softer $D_{F}(2r)$ in the bilocal theory. Hence the interpretation of the NJL $M^{2}$ as a loop correction to the mass is now replaced in the bilocal theory as a loop correction to the potential, enhancing binding.
Hence, in the present analysis we will use the NJL limit in obtaining approximate bilocal loop integral results, but judiciously replacing $\delta^{4}(x-y)\rightarrow M_{0}^{2}D_{F}(x-y)$. Physically, the bilocal theory has a true momentum space cutoff of loop integrals $\sim M_{0}^{2}$, due to $D_{F}(q^{2})$, while in the NJL theory this cut-off is implemented in loop integrals “by hand.” The replacement of the $\delta^{4}(x-y)$ by $M_{0}^{2}D_{F}(x-y)$ approximately implements the true physics of the bilocal theory, allowing us to obtain a potential that, in turn, allows the “back-reaction” of the solution. We are invoking the “indistinguishability” of the $\delta^{4}(x-y)$ from the bilocal $M_{0}^{2}D_{F}(x-y)$, but maintaining the bilocality of the present theory. By extending the auxiliary field equation to include the additional terms of $W(\Omega)$, then integrating out $B_{i}$, we obtain the quantum loop corrections to the semiclassical action. These are the analogues of the terms in eqs.(64,65) generated by eq.(63), but now for the bilocal fields.
To compute $W(\Omega)$ we compute the path integral of eq.(II.1), integrating out $\Omega$. The coupling of $B_{i}$ to the underlying $\omega$-fields is implicit in eqs.(10,II.1):
$\int_{xy}\biggl{(}g_{0}\overline{\omega}_{R}(y)D_{F}(x-y)\omega_{iL}(x)B^{i\dagger}(x,y)+h.c.\biggr{)}.$ (66)
This yields single-particle irreducible (1PI) diagrams leading to quadratic and quartic terms in $B$:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{i}{\hbar}W(\Omega)=\!\int_{xyx^{\prime}y^{\prime}}\!\!\!\![B^{\dagger}(x,y){\cal{F}}(x,y,x^{\prime},y^{\prime})B(x^{\prime},y^{\prime})],$
$+\!\int_{x...z^{\prime}}\!\!\!\![B^{\dagger}(x,y)B(x^{\prime},y^{\prime})]{\cal{G}}(x,y,x^{\prime},y^{\prime},w,z,w^{\prime},z^{\prime})[B^{\dagger}(w,z)B(w^{\prime},z^{\prime})]+$... (67)
which largely parallels the loop corrections of the NJL model. We give the final results of the loop calculations below in eqs.(III.3,III.3).
III.1 Main Loop
Note the first term of eq.(III) contains the 4-point function ${\cal{F}}(x,y,x^{\prime},y^{\prime})$, given by:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{i}{\hbar}W_{\cal{F}}(\Omega)=\!\int_{xyx^{\prime}y^{\prime}}\!\![B^{\dagger}(x,y){\cal{F}}(x,y,x^{\prime},y^{\prime})B(x^{\prime},y^{\prime})]=$
$-\frac{1}{2}g_{0}^{2}N_{c}\!\int_{x...y^{\prime}}\!\!\![B^{\dagger}(x,y)D_{F}(x-y)\mathop{\rm Tr}(S_{F}(x-x^{\prime})S_{F}(y^{\prime}-y))D_{F}(x^{\prime}-y^{\prime})B(x^{\prime},y^{\prime})],$ (68)
where $S_{F}(x)$ is a fermionic Feynman propagator (we use conventions of BjDrell ):
$S_{F}(x)=\int\frac{d^{4}\ell}{(2\pi)^{4}}\frac{i\ell\!\!\!/\!\,\,}{\ell^{2}+i\epsilon}e^{i\ell\cdot x}\qquad\makebox{and}\qquad\mathop{\rm Tr}\biggl{(}\frac{(1-\gamma^{5})}{2}\ell\!\!\!/\!\,\,\;\ell\!\!\!/\!\,\,^{\prime}\biggr{)}=2\ell\cdot\ell^{\prime}.$ (69)
In the loop integral we make the NJL approximation of taking a pointlike limit of the integral, defined by:
$D_{F}(x-y)=-\int\frac{d^{4}q}{(2\pi)^{4}}\frac{e^{iq(x-y)}}{q^{2}-M_{0}^{2}}\rightarrow\int\frac{d^{4}q}{(2\pi)^{4}}\frac{e^{iq(x-y)}}{M_{0}^{2}}=\frac{1}{M_{0}^{2}}\delta^{4}(x-y),$ (70)
then,
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!{\cal{F}}(x,y,x^{\prime},y^{\prime})\rightarrow-2g_{0}^{2}N_{c}\frac{\delta^{4}(x-y)\delta^{4}(x^{\prime}-y^{\prime})}{M_{0}^{4}}\int\frac{d^{4}\ell}{(2\pi)^{4}}\frac{d^{4}\ell^{\prime}}{(2\pi)^{4}}\frac{\ell\cdot\ell^{\prime}e^{i(\ell-\ell^{\prime})(x-x^{\prime})}}{\ell^{2}\ell^{\prime 2}}.$ (71)
It is useful to go to barycentric coordinates:
$X=\frac{1}{2}(x+y),\;\;\;r=\frac{1}{2}(x-y),\;\;\;J=\left|\frac{\partial(x,y)}{\partial(X,r)}\right|=2^{4},\;\;\;$ (72)
and define:
$\widetilde{B}(X,r)\equiv B(X-r,X+r).$ (73)
We will also define new integration variables, $\overline{X}=(X+X^{\prime})/2$ and $R=(X-X^{\prime})$, which have unit Jacobian.
Hence we have, with $\delta^{4}(x-y)=J\delta^{4}(2r)$,
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\int_{xyx^{\prime}y^{\prime}}\![B^{\dagger}(x,y){\cal{F}}(x,y,x^{\prime},y^{\prime})B(x^{\prime},y^{\prime})]$
$\!\!\!\!\!\!\!\!\!\!\rightarrow-g_{0}^{2}N_{c}J^{2}\!\int_{XrX^{\prime}r^{\prime}}\!\!\!\!\![\widetilde{B}^{\dagger}(X,r)\frac{\delta^{4}(2r)\delta^{4}(2r^{\prime})}{M_{0}^{4}}\int\frac{d^{4}\ell}{(2\pi)^{4}}\frac{d^{4}\ell^{\prime}}{(2\pi)^{4}}\frac{2\ell\cdot\ell^{\prime}e^{2i(\ell-\ell^{\prime})(X+r-X^{\prime}-r^{\prime})}}{\ell^{2}\ell^{\prime 2}}\widetilde{B}(X^{\prime},r^{\prime})]$
$\!\!\!\!\!\!\!\!\!\!=-\frac{g_{0}^{2}N_{c}}{M_{0}^{4}}\!\int_{XX^{\prime}}\![\widetilde{B}^{\dagger}(X,0)\int\frac{d^{4}\ell}{(2\pi)^{4}}\frac{d^{4}\ell^{\prime}}{(2\pi)^{4}}\frac{2\ell\cdot\ell^{\prime}}{\ell^{2}\ell^{\prime 2}}e^{i(\ell-\ell^{\prime})(X-X^{\prime})}\widetilde{B}(X^{\prime},0)].$ (74)
Note the cancellation of $J^{2}=2^{8}$ with the $(1/2^{4})^{2}$ coming from integrating $\delta^{4}(2r)\delta^{4}(2r^{\prime})$. Then $\widetilde{B}^{\dagger}(X,0)$ ($\widetilde{B}(X^{\prime},0)$) is reduced to a pointlike field with dependence only upon $X$ ($X^{\prime}$) where $r,r^{\prime}\rightarrow 0$.
An outgoing (incoming) state is, $\widetilde{B}^{\dagger}(X,0)\sim\widetilde{B}^{\dagger}_{0}e^{iP(\overline{X}+R/2)}$, ( $\widetilde{B}(X^{\prime},0)\sim\widetilde{B}_{0}e^{-iP(\overline{X}-R/2)}$). Thus, integrating over $(\overline{X},R)$ we obtain,
$-\frac{g_{0}^{2}N_{c}}{M_{0}^{4}}\!\int_{{X}R}e^{i(\ell-\ell^{\prime}+P)(R)}[\widetilde{B}^{\dagger}_{0}{}^{\prime}\widetilde{B}_{0}]\int\frac{d^{4}\ell}{(2\pi)^{4}}\frac{2\ell\cdot(\ell^{\prime})}{\ell^{2}(\ell^{\prime})^{2}}=-\frac{2g_{0}^{2}N_{c}}{M_{0}^{4}}\!\int_{{X}}[\widetilde{B}^{\dagger}_{0}{}^{\prime}\widetilde{B}_{0}]\int\frac{d^{4}\ell}{(2\pi)^{4}}\frac{\ell\cdot(\ell+P)}{\ell^{2}(\ell+P)^{2}}$
$\qquad\qquad=\frac{ig_{0}^{2}N_{c}}{8\pi^{2}M_{0}^{4}}\!\int_{\overline{X}}[\widetilde{B}_{0}^{\dagger}\widetilde{B}_{0}]\biggl{(}M_{0}^{2}+P^{2}\ln\left(\frac{M_{0}}{m}\right)\biggr{)}.$ (75)
Here we use a Wick rotation and Euclidean momentum UV cut-off, $\ell^{2}<M_{0}^{2}$.
We have obtained the analogues of the $M^{2}$ and $P^{2}$ terms in eqs.(64,65) acting on the auxiliary field. The $P^{2}$ term corresponds to a loop-generated kinetic term for the auxiliary field $B_{i}$, (see eq.(94)). In the framework of the NJL model and original top condensation theory, BHL , this corresponds the formation of a physical state, i.e., a resonance. We are interested presently in small $P^{2}$ so we will treat the $P^{2}\approx 0$ limit of the loop and return to this issue below.
The remaining $M_{0}^{2}$ term of eq.(III.1) is just the quadratically divergent two-point loop of the NJL model in disguise, as in eq.(65). We multiply by $i\hbar$ to obtain the two-point loop contribution to the action $W_{\cal{F}}(\Omega)$,
$W_{\cal{F}}(\Omega)=\frac{g_{0}^{2}N_{c}}{8\pi^{2}M_{0}^{2}}\!\int d^{4}X\;[\widetilde{B}^{\dagger}(X,0)\widetilde{B}(X,0)]\rightarrow\frac{g_{0}^{2}N_{c}J}{8\pi^{2}M_{0}^{2}}\!\int d^{4}Xd^{4}r\;\widetilde{B}^{\dagger}(X,r)\delta^{4}(2r)\widetilde{B}(X,r).$ (76)
Note the positive sign in the action, corresponding to an attractive potential. While in the pointlike NJL model this term was interpreted as the quadratic negative term in the potential, $g^{2}N_{c}M_{0}^{2}/8\pi^{2}$, we see that, in the bilocal theory due to the implicit $\delta^{4}(2r)$, this is a potential and not a mass term. This presently acts as an attractive $\delta^{4}(2r)$ potential, and the wave-function $\phi(r)$ will react to it as such. However, as stated above, the $\delta^{4}(2r)$ is indistinguishable from the tree-level potential via: $\delta^{4}(2r)\rightarrow M_{0}^{2}D_{F}(x-y).$ Therefore, using this replacement of the pointlike $\delta$-function we obtain in the action $(P^{2}\approx 0)$:
$W_{\cal{F}}(\Omega)=\!\int_{x...y^{\prime}}\!\!\!B^{\dagger}(x,y){\cal{F}}(x,y,x^{\prime},y^{\prime})B(x^{\prime},y^{\prime})\approx\frac{g_{0}^{2}N_{c}}{8\pi^{2}}\int d^{4}x\;d^{4}y\;B^{\dagger}(x,y)D_{F}(x-y)B(x,y),$ (77)
where we have returned to the original $(x,y)$ coordinates, and we have a two-body interaction. This is loop level, $\cal{O}$$\hbar$, and upon integrating out the auxiliary field, $B$, this will yield an enhancement of the semiclassical potential.
This result answers the potential criticism of main , mentioned in the Introduction: If one computes the loop correction to the BEH mass by naively looping the Yukawa interaction in the semiclassical theory, then one would obtain the result of eq.(III.1), which appears to be a large correction to the BEH mass $\propto M_{0}^{2}$. Of course, this would be falsely double counting, since the only occurrence of the loop is in the underlying theory. Moreover, as emphasized above, this is not a mass term for the bilocal field since it is $\propto\ \delta^{4}(2r)$ and, rather, it is a short-distance (attractive) correction to the potential for $\phi(r)$. The $\phi(r)$ field will adjust accordingly as the solution to the SKG equation, multiplicatively modifying the eigenvalue $\mu^{2}$. Indeed, due to effective large distance scale symmetry near critical coupling, this must be interpreted as a radiative (multiplicative) correction to the effective coupling in the semiclassical potential. The loop integral leads to an attractive interaction which represents an enhancement of the effective coupling. We elaborate this further below.
III.2 Quartic Coupling
We also obtain the induced quartic interaction, which appears as an 8-point function in the effective semiclassical theory. This can be likewise evaluated in the NJL approximation. The loop integral involves $D_{F}^{4}(2r)\mathop{\rm Tr}{(S_{F})^{4}}$ and we replace three of the $D_{F}$ factors with $\delta^{4}(2r)/M_{0}^{2}$ which leaves one two-point integral of order $g_{0}^{4}$. The result can be inferred from the NJL calculation of the $\lambda$ term in eq.(65) and is consistent with the detailed calculation:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!W_{\cal{G}}(\Omega)=\int_{x...z^{\prime}}\!\!\!\!\!\!\!\!B^{\dagger}(x,y)B(x^{\prime},y^{\prime}){\cal{G}}(x,y,x^{\prime},y^{\prime},w,z,w^{\prime},z^{\prime})B^{\dagger}(w,z)B(w^{\prime},z^{\prime})$
$\qquad\approx-\frac{{g}^{4}_{0}\hat{\lambda}}{2M_{0}^{8}}\biggl{(}\!\int_{xy}\!\!({B}^{\dagger}(x,y){B}(x,y))^{2}\delta^{4}(x-y)\biggr{)},$ (78)
where we define:
$\;\;\hat{\lambda}=\frac{N_{c}}{4\pi^{2}}\ln\biggl{(}\frac{M_{0}}{m}\biggr{)}.$ (79)
We develop this further below.
III.3 Full Renormalized Action
The resulting action with eq.(44) and the principal quantum effects considered here is then :
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S=S_{KTH}+W(\Omega)+\!\int_{x}\biggl{(}[\overline{\psi}_{L}iD\!\!\!/\!\,\,\psi_{L}]_{f}+[\overline{\psi}_{R}\;iD\!\!\!/\!\,\,\psi_{R}]_{f}\biggr{)}+\int_{xy}\;\biggl{(}-[B^{i}{}^{\dagger}D_{F}B_{i}]+g_{0}(B^{i\dagger}D_{F}C_{i}+h.c.)\biggr{)}.$ (80)
We recall eq.(30), $C_{i}(x,y)=\sqrt{N_{c}}M_{0}^{2}H_{i}(x,y)+[\overline{\psi}_{R}(y)\psi_{iL}(x)]_{f}$, and we have from the loop calculations:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!W(\Omega)=\!\int_{xyx^{\prime}y^{\prime}}\!\!\!\![B^{\dagger}(x,y){\cal{F}}(x,y,x^{\prime},y^{\prime})B(x^{\prime},y^{\prime})]+\!\int_{x...z^{\prime}}\!\!\!\![B^{\dagger}(x,y)B(x^{\prime},y^{\prime})]{\cal{G}}(x,...,z^{\prime})[B^{\dagger}(w,z)B(w^{\prime},z^{\prime})]+$...,
$=\frac{g_{0}^{2}N_{c}}{8\pi^{2}}\int_{xy}\;B^{\dagger}(x,y)D_{F}(x-y)B(x,y)-\frac{g_{0}^{4}\hat{\lambda}}{2M_{0}^{8}}\!\int_{xy}\!\!({B}^{\dagger}(x,y){B}(x,y))^{2}\delta^{4}(x-y).$ (81)
Hence $S$ can be can be rewritten as:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S=S_{KTH}+\!\int_{x}\biggl{(}[\overline{\psi}_{R}iD\!\!\!/\!\,\,\psi_{R}]_{f}+[\overline{\psi}_{L}iD\!\!\!/\!\,\,\psi_{L}]_{f}\biggr{)}+\!\int_{xy}\!\biggl{(}g_{0}B^{i\dagger}(x,y)D_{F}(x-y)C_{i}+h.c.\biggr{)}$
$-\!\int_{xy}\![B^{\dagger}(x,y)D_{F}(x-y)B(x,y)]\biggl{(}1-\frac{g_{0}^{2}N_{c}}{8\pi^{2}}\biggr{)}-\frac{g_{0}^{4}\hat{\lambda}}{2M_{0}^{8}}\int_{xy}\![B^{\dagger}(x,y)B(x,y)]^{2}\delta^{4}(x-y).$ (82)
We now rescale the auxiliary field $B_{i}$ and the coupling $g_{0}$:
$B^{\prime}_{i}=\sqrt{R}B_{i}(x,y)\qquad\overline{g}_{0}=R^{-1/2}g_{0}\qquad\text{where,}\qquad R=\biggl{(}1-\frac{g_{0}^{2}N_{c}}{8\pi^{2}}\biggr{)}.$ (83)
We then have,
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S=S_{KTH}+\!\int_{x}\biggl{(}[\overline{\psi}_{R}iD\!\!\!/\!\,\,\psi_{R}]_{f}+[\overline{\psi}_{L}iD\!\!\!/\!\,\,\psi_{L}]_{f}\biggr{)}+\!\int_{xy}\!\biggl{(}\overline{g}_{0}B^{\prime}_{i}{}^{\dagger}(x,y)D_{F}(x-y)C^{i}(x,y)+h.c.\biggr{)}$
$-\!\int_{xy}\![B^{\prime\dagger}(x,y)D_{F}(x-y)B^{\prime}(x,y)]-\frac{\overline{g}_{0}^{4}\hat{\lambda}}{2M_{0}^{6}}\int_{xy}\!D_{F}(x-y)[B^{\prime\dagger}(x,y)B^{\prime}(x,y)]^{2}+$..., (84)
where in the last term we use $\delta^{4}(x-y)\approx M_{0}^{2}D_{F}(x-y)$.
The “equation of motion” of the auxiliary field, $B^{\prime}$, using eq.31, then takes the form,
$B^{\prime}_{i}(x,y)=\overline{g}_{0}C_{i}(x,y)-\overline{g}_{0}^{4}\hat{\lambda}M_{0}^{-6}B^{\prime}_{i}(x,y)(B^{\prime}{}^{j}{}^{\dagger}(x,y)B^{\prime}_{j}(x,y))+$..., (85)
where we defined $C_{i}$ in eq.(30). We then solve eq.(85) perturbatively in $\hat{\lambda}$ (where $\hat{\lambda}$ is defined in eq.(79):
$B^{\prime}_{i}(x,y)\approx\overline{g}_{0}C_{i}-\overline{g}^{7}_{0}\hat{\lambda}M_{0}^{-6}C_{i}(C^{\dagger j}(x)C_{j}(x))+$... (86)
Substituting into eq.(III.3), and noting cancellations (particularly leading to the quartic term), the action is then:
$S=S_{KTH}+\!\int_{x}\!\biggl{(}[\overline{\psi}_{L}(x)iD\!\!\!/\!\,\,\psi_{L}(x)]+[\overline{\psi}_{R}(x)\;iD\!\!\!/\!\,\,\psi_{R}(x)]\biggr{)}$
$+\overline{g}_{0}^{2}\!\int_{xy}\biggl{(}N_{c}M_{0}^{4}H^{\dagger}(x,y)D_{F}(x-y)H(x,y)+\;\sqrt{N_{c}}M_{0}^{2}(H^{i\dagger}(x,y)D_{F}(x-y)[\overline{\psi}_{R}(y)\psi_{iL}(x)]_{f}+h.c.)$
$-[\overline{\psi}^{i}_{L}(x)\psi_{R}(y)]_{f}D_{F}(x-y)[\overline{\psi}_{R}(y)\psi_{iL}(x)]_{f}\biggr{)}-\frac{1}{2}\overline{g}^{8}_{0}N_{c}\hat{\lambda}\int_{xy}\!\delta^{4}(x-y)[H^{\dagger i}(x,y)H_{i}(x,y)]^{2}+$... (87)
Therefore, we see that, having integrated out the auxiliary field, we have obtained an enhanced coupling in the $0^{+}$ binding channel:
$\overline{g}_{0}^{2}={g_{0}^{2}}\left(1-\frac{N_{c}g_{0}^{2}}{8\pi^{2}}\right)^{-1}.$ (88)
The rescaling of $g_{0}$ summed the tower of fermion loops $\propto(g_{0}^{2}N_{c}/8\pi^{2})^{n}$, which is seen by expanding $R^{-1}$. The renormalized coupling $\overline{g}_{0}$ applies to the binding action, and not to the interaction of free fermions with the colorons (e.g., for the decay width of a coloron we would use the unrenormalized $g_{0}$). On energy scales $M_{0}$ the enhanced $\overline{g}_{0}$ applies in the 4-fermion interaction of free fermions, as is indicated in the effective action in eq.(III.3). However, for energy scales $M_{0}$ the theory reverts to a pure topcolor gauge theory with coupling $g_{0}$ and the binding effects disappear. We see that $\overline{g}_{0}>g_{0}$, so the underlying topcolor theory is weaker than the effective binding interaction. This is an example of a “critical amplification” effect in the binding theory.
Passing to barycentric coordinates, $H_{i}(x,y)\rightarrow\sqrt{2/J}\;H_{i}(X)\phi(r),$ and, $M_{0}^{2}D_{F}(2r)\rightarrow\delta^{4}(2r)$, we see that the Yukawa interaction becomes:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\overline{g}_{0}^{2}\!\int_{xy}\!\!\!\!\sqrt{N_{c}}M_{0}^{2}(H^{i\dagger}(x,y)D_{F}(x-y)[\overline{\psi}_{R}(y)\psi_{iL}(x)]_{f}+h.c.)\approx\overline{g}_{Y}\int d^{4}X\biggl{(}H^{i\dagger}(X)[\overline{\psi}_{R}(X)\psi_{iL}(X)]_{f}+h.c.\biggr{)}.$ (89)
Therefore, we now have a renormalized Yukawa coupling:
$\overline{g}_{Y}=\overline{g}_{0}^{2}\sqrt{2N_{c}/J}\;\phi(0).$ (90)
The physical top-Yukawa coupling is now given by $\overline{g}_{Y}$ and this will be set to its experimental value of unity as an input to define the theory and obtain $M_{0}$.
Furthermore, we have for the quartic term (where $\hat{\lambda}$ is defined in eq.(79):
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!-\frac{1}{2}\overline{g}^{8}_{0}N_{c}\hat{\lambda}\int_{xy}\!\delta^{4}(x-y)[H^{\dagger i}(x,y)H_{i}(x,y)]^{2}\longrightarrow-\frac{1}{2}\hat{\lambda}[\overline{g}^{8}_{0}JJ^{-1}(2/J)^{2}N_{c}^{2}|\phi(0)|^{4}]\int_{X}(H^{i\dagger}(X)H_{i}(X))^{2}$
[…]
We can therefore write the leading terms in the full theory:
$S=\!\int_{x}\!\biggl{(}[\overline{\psi}_{L}(x)iD\!\!\!/\!\,\,\psi_{L}(x)]_{f}+[\overline{\psi}_{R}(x)\;iD\!\!\!/\!\,\,\psi_{R}(x)]_{f}\biggr{)}$
$+\!\int_{X}\biggl{(}D_{\mu}H^{\dagger}(X)D^{\mu}H(X)+H^{\dagger}(X)H(X)\int_{r}\biggl{(}-|\nabla_{\vec{r}}\phi(\vec{r})|^{2}+\left(\overline{g}_{0}^{2}D_{F}(2r)|\phi(r)|^{2}\right)\biggr{)}$
$\qquad+\overline{g}_{Y}(H^{i\dagger}(X)[\overline{\psi}_{R}(X)\psi_{iL}(X)]_{f}+h.c.)-\frac{\lambda}{2}\int_{X}(H^{\dagger}(X)H(X))^{2}\biggr{)}$
$+\;\overline{g}_{0}^{2}\!\int_{xy}\;[\overline{\psi}^{i}_{L}(x)\psi_{R}(y)]_{f}D_{F}(x-y)[\overline{\psi}_{R}(y)\psi_{iL}(x)]_{f}+$... (92)
With a solution to the SKG equation for $\phi(r)$ with eigenvalue $\mu^{2}$ (and relabeling the now dummy integration variable $X\rightarrow x$) we have:
$S=\!\int_{x}\!\biggl{(}[\overline{\psi}_{L}(x)iD\!\!\!/\!\,\,\psi_{L}(x)]_{f}+[\overline{\psi}_{R}(x)\;iD\!\!\!/\!\,\,\psi_{R}(x)]_{f}\biggr{)}$
$+\!\int_{x}\biggl{(}D_{\mu}H^{\dagger}(x)D^{\mu}H(x)-\mu^{2}H^{\dagger}(x)H(x)+\;\overline{g}_{Y}(H^{i\dagger}(x)[\overline{\psi}_{R}(x)\psi_{iL}(x)]_{f}+h.c.)-\frac{\lambda}{2}(H^{\dagger}(x)H(x))^{2}\biggr{)}$
$+\;\overline{g}_{0}^{2}\!\int_{xy}\;[\overline{\psi}^{i}_{L}(x)\psi_{R}(y)]_{f}D_{F}(x-y)[\overline{\psi}_{R}(y)\psi_{iL}(x)]_{f}+$... (93)
where $\mu^{2}<0$ for supercritical $\overline{g}_{0}^{2}$. We have recovered the Standard Model BEH action with couplings to third generation $(t,b)$ quarks and the novel 4-fermion interaction associated with coloron exchange. The covariant derivative, $D_{\mu}$, is that of the BEH boson eq.(41) (viewed either as a low energy limit or explicitly with the incorporation of Wilson lines as discussed in main ).
While we have extrapolated the loops to zero $P^{2}$, we note that eq.(III.1) contains a $P^{2}$ term. If we interpret this in the Wilsonian description of the NJL model, as in main ; BHL , then eq.(III.1) implies that the theory is creating a loop level resonance of the $B$ field with an action (ignoring gauge fields):
$\sim\widetilde{Z}\partial_{X}B^{\prime i\dagger}\partial_{X}B^{\prime}_{i}-M_{0}^{2}B^{\prime i\dagger}B^{\prime}_{i}\qquad\text{where,}\qquad\widetilde{Z}=\frac{\overline{g}_{0}^{2}N_{c}}{8\pi^{2}}\ln\left(\frac{M_{0}}{m}\right),$ (94)
where $m\sim|P|$. The on-shell resonance mass near critical coupling is therefore $M_{resonance}^{2}=M_{0}^{2}/\ln\left(\frac{M_{0}}{m}\right).$ (Here we have reverted to interpreting $M_{0}^{2}D_{F}\rightarrow\delta^{4}(2r)$). Then eq.(86) is satisfied as $P^{2}<\!\!<M_{0}^{2}$. The formation and exchange of this resonance below threshold $P^{2}\sim 0$ generates an attractive interaction and can be viewed as leading to the enhancement of the effective coupling $\overline{g}_{0}^{2}$ as $P^{2}\rightarrow 0$.
IV Discussion and Conclusions
IV.1 Physical Consequences of the Improved Yukawa Coupling
The renormalized topcolor coupling in the $0^{+}$ binding channel satisfies:
$\overline{g}_{0}^{2}={g_{0}^{2}}\biggl{(}1-\frac{N_{c}g_{0}^{2}}{8\pi^{2}}\biggr{)}^{-1}.$ (95)
where $g_{0}^{2}$ is the underlying topcolor coupling. This represents “critical amplification” of the underlying coupling $g_{0}^{2}$ and all of the induced couplings, $g_{Y}$ and $\lambda$ (etc.). It is $\overline{g}_{0}^{2}$ that controls the formation of the semiclassical bound state. This amplification effect is not unexpected, since the quantum fermion loops are essentially NJL model effects, and these are strongly binding. We see that this can be understood as the formation of a resonance in the $0^{+}$ as an NJL bound state, enhancing the effective coupling $\overline{g}_{0}^{2}$ as $P^{2}\rightarrow 0$.
The bound state is critical (has zero mass $\mu^{2}$) when $\overline{g}_{0}^{2}=\overline{g}_{0c}^{2}$ where we determine numerically main :
$\frac{\overline{g}_{0c}^{2}N_{c}}{8\pi^{2}}=1.06940,\qquad\text{cf. NJL model:}\;\;\frac{\overline{g}_{0c}^{2}N_{c}}{8\pi^{2}}=1.$ (96)
This is the criticality condition of the Schrödinger equation with a Yukawa potential as translated into the SKG equation (note that we’ve often approximated $g_{0c}^{2}\approx 8\pi^{2}/N_{c}$ using the NJL critical value in main ).
Solutions to the SKG equation lead to the eigenvalue $\mu^{2}$ which can be identified with the SM symmetric phase, $\mu^{2}=-(88)^{2}$ GeV2, which is negative (tachyonic) for supercritical $\overline{g}_{0}^{2}>\overline{g}_{0c}^{2}$. The solutions are discussed in main . For example, if we assume a large distance solution of the form (the “skeletal solution” of Section IV.E of main ), then
$\phi(r)\approx\frac{ce^{-|\mu|r}}{r};\;\;\;\;\;\makebox{then we find,}\;\;\;\;\;\phi(0)=\frac{2\sqrt{2}}{\pi^{3/2}}\sqrt{\frac{|\mu|}{M_{0}}}=0.50795\sqrt{\frac{|\mu|}{M_{0}}}.$ (97)
We can then fit the renormalized Yukawa coupling to experiment,
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\overline{g}_{Y}=1=\overline{g}^{2}_{0c}\sqrt{\frac{2N_{c}}{J}}\;\phi(0)=17.236\;\phi(0)\approx 8.7548\sqrt{\frac{|\mu|}{M_{0}}},$ (98)
we then predict $M_{0}=6.745\;\makebox{TeV}$. [^8] This was computed in main where we quoted $5.9$ TeV, which resulted from using the NJL critical value $N_{c}g_{0c}^{2}/8\pi^{2}=1$ rather than the true $N_{c}g_{0c}^{2}/8\pi^{2}=1.0694$. However, we think these are within the uncertainty of the leading order calculations for the SKG solutions, and does not take RG running effects, e.g., logarithms, into account. Likewise, using an approximate numerical solution to the SKG equation yields $M_{0}=5.23\;\makebox{TeV}$ (this is sensitive to a fitting function, and likely underestimates $M_{0}$). Conservatively we would estimate $M_{0}$ to lie in a range of order $5<M_{0}<7$ TeV.
It should be emphasized that the collider physics predictions for properties of colorons will depend mainly upon $g_{0}^{2}$, not $\overline{g}_{0}^{2}$. The colorons form a QCD color octet, but with the exception of couplings to $\overline{t}t$ would not be singly produced and would be pair produced in gluon fusion processes, which is sensitive to the QCD coupling and $M_{0}$. Such coloron pairs lie near the energy limit of the LHC and will decay to pairs of $\overline{t}t$, and we also expect $\overline{b}b$ (topcolor must be extended to include the $b_{R}$-quark and a $Z^{\prime}$ to insure the $\overline{b}b$ channel is subcritical Topcolor ). Coloron decay widths are controlled by $g_{0}^{2}$. If $\overline{g}_{0}^{2}$ is approximately critical, $N_{c}\overline{g}_{0}^{2}/8\pi^{2}\approx(1.0694)$ then from eq.(95) we have,
$\frac{g_{0}^{2}N_{c}}{8\pi^{2}}\approx\frac{1.0694}{1+1.0694}=0.51677,$ (99)
and the underlying topcolor theory is subcritical, and $\alpha_{topcolor}=\frac{g_{0}^{2}}{4\pi}\approx 1.0823$ is marginally perturbative.
IV.2 Physical Consequences of the Improved Quartic Coupling
For the quartic coupling we obtain above:
$\!\!\!\!\!\!\!\!\!\!{\lambda}\approx\frac{\overline{g}_{Y}^{4}N_{c}}{4\pi^{2}}\ln\left(\frac{M_{0}}{\mu}\right).$ (100)
The log evolution is just that of a single fermion loop and matches the result for the pointlike NJL case, with $g_{top}=\overline{g}_{Y}$. Not surprisingly when we take the pointlike potential limit the loop result of the bilocal theory confirms a pointlike NJL calculation (see discussion of Section 4 of ref.(main2 ).
Experimentally, in the SM using the value of $m_{BEH}\approx 125$ GeV and $v_{weak}\approx 175$ GeV, we find $\lambda\approx 0.25$. In the old pointlike NJL top condensation model the quartic coupling was determined by the RG with “compositeness boundary conditions,” where we obtained (running-down from the Landau pole at $M_{0}$ to $|\mu|$), the result $\lambda\sim 1$. This is too large and leads to predicted $m_{BEH}\sim 260$ GeV. Indeed, the quartic coupling is generally problematic for NJL based theories of a composite BEH boson.
However, in the present bilocal scheme owing to suppression of $\overline{g}_{Y}\propto\phi(0)$, the quartic coupling is also suppressed $\propto|\phi(0)|^{4}$ and is now generated in RG running, from a value of zero at $M_{0}=6$ TeV, down to $|\mu|\sim 88$ GeV, using the induced $\overline{g}_{Y}$. Keeping only the $\overline{g}_{Y}^{4}$ contribution we would obtain numerically from eq.(100):
$\lambda\approx(\overline{g}_{Y}^{4})\frac{N_{c}}{4\pi^{2}}\ln\left(\frac{M_{0}}{|\mu|}\right)\approx 0.321\;\;\;\makebox{(cf., $0.25$ experiment).}$ (101)
This result is significantly better than the old NJL top condensation model. However, prefactor at one loop level should reflect the full RG running of $\lambda$, (see, e.g., HillTA ), which would yield at leading log approximation:
$\lambda\approx(\overline{g}_{Y}^{4}-\overline{g}_{Y}^{2}\lambda-\lambda^{2})\frac{N_{c}}{4\pi^{2}}\ln\left(\frac{M_{0}}{|\mu|}\right)\approx 0.230\;\;\;\makebox{(cf., $0.25$ experiment).}$ (102)
This includes the $\lambda^{2}$ term which arises from loops with propagating internal BEH boson. For the present composite model this not likely to be correct, due to the extended size of the BEH wave-function. The terms of order $\overline{g}_{Y}^{2}\lambda$ are leg-renormalizations, due to the top quark loop and presumably can be retained. This would imply a result:
$\lambda\approx(\overline{g}_{Y}^{4}-\overline{g}_{Y}^{2}\lambda)\frac{N_{c}}{4\pi^{2}}\ln\left(\frac{M_{0}}{\mu}\right)\approx 0.243\;\;\;\makebox{(cf., $0.25$ experiment).}$ (103)
IV.3 Summary
We have used the techniques of Jackiw, et al. ,\!\!, Jackiw ; CJT to obtain formally the effective action for the semiclassical third generation fields $\psi_{L},\psi_{R}$ in a single coloron exchange interaction field theory of topcolor. This describes the Brout-Englert-Higgs boson as a composite object with constituents that are the semiclassical fields composed of $t$ and $b$ quarks.
This marks a radical departure from the old RG improved NJL model which builds a composite state by integrating out all of the fermions, in analogy to confinement in QCD. For us, the bound state is softer, and more Hydrogenic, containing a subset of the low momentum modes in the semiclassical theory. This gives us the internal wave-function $\phi(r)$. Tuning the theory within a few percent of the critical coupling yields a hierarchy that defines a low mass BEH bound state. We have given the central results in the companion paper main and addressed the most significant quantum effects presently. Here we have introduced an auxiliary field, $B_{i}$, which consolidates the main large-$N_{c}$ fermion loop contributions. Upon computing these, we integrate out $B_{i}$ to obtain the full quantum corrected action.
Our most significant new result is that the topcolor coupling, $g_{0}^{2}$ is subcritical while the effective coupling in the binding potential $\overline{g}_{0}^{2}$ is enhanced and can be slightly supercritical:
$\overline{g}_{0}^{2}={g_{0}^{2}}\biggl{(}{1-\frac{N_{c}g_{0}^{2}}{8\pi^{2}}}\biggr{)}^{-1}$ (104)
This in turn renormalizes the Yukawa coupling and quartic coupling. Hence the value of $\overline{g}_{0}^{2}$ is “critically amplified” above a weaker value of the underlying topcolor coupling $g_{0}^{2}$.
Inputting the experimental result for the renormalized Yukawa coupling, $\overline{g}_{Y}=1$ with the symmetric phase BEH mass, $\mu^{2}=-(88)^{2}$ GeV2, yields the central prediction obtained for $M_{0}\approx 6$ TeV as discussed in main . Hence the major results of main still hold, but with $g_{0}^{2}$ now replaced by $\overline{g}_{0}^{2}$. Due to the linear relationship between $\overline{g}_{0}^{2}$ and $|\mu|/M_{0}$, a consequence of the dilution effect of $\phi(0)$ as discussed in main , we see the degree of fine-tuning of the hierarchy is of order $\sim|\mu|/M_{0}\sim 1.4\%$.
We explicitly demonstrate that the BEH boson quartic coupling, $\lambda$, arises from loops $\sim N_{c}\overline{g}_{Y}^{4}\ln(M_{0}/\mu)/4\pi^{2}\propto(\phi(0))^{4}$. The result for the quartic coupling is $\lambda\sim 0.24$, compared to $\lambda\approx 0.25$ experimentally. The $\lambda$ prediction is in astonishingly good agreement with experiment and we believe makes the natural top condensation idea compelling.
The “colorons” which mediate the binding interaction form a QCD color octet. They have a global $SU(3)$ symmetry and a conserved $SU(3)$ current, and must therefore be pair-produced. They will decay in the minimal model to $\overline{t}t$, but in more realistic topcolor models they can also decay to $\overline{b}{b}$ (Topcolor Topcolor2 NSD and references therein). They may be accessible to the LHC in the multi-TeV range Burdman ; expt ; limits , favoring the third generation in its couplings. It should be straightforward to obtain lower bounds on $M_{0}$ from single gluon or gluon fusion production of pairs of the colorons. The colorons will produce excesses in 4-top, $t\overline{t}t\overline{t}$, events. In an extended model we would also expect $t\overline{t}b\overline{b}$ and $b\overline{b}b\overline{b}$ anomalies to emerge. The more general topcolor models offer many possibilities within the arena of heavy quark flavor physics.
This theory, if confirmed, solves the “naturalness problem” of the BEH boson in the Standard Model and opens up a vista of new physics with new gauge interactions. Many avenues for further theoretical development exist. Notably, a revisitation of the Topcolor $Z^{\prime}$ Topcolor , resonances, and new flavor physics associated with $\overline{b}b$ would be of interest. There is much to do to further develop these models in the context of the present dynamics.
Acknowledgments
I thank Bill Bardeen, Thom Curtright, Bogdan Dobrescu, Andreas Kronfeld and Julius Kuti for discussions and comments during the course of this work.